OpenStudy (anonymous):

Write the inequality in slope-intercept form: 9x+3y=>-12

OpenStudy (anonymous):

\[9x+3y \ge-12\]

OpenStudy (anonymous):

slope intercept formk is y=mx+b

OpenStudy (anonymous):

that doesnt help me...

OpenStudy (anonymous):

put that into y=mx+b and there is ur answer sorry but I cant help u any more than that idk how to do it i just kno that y=mx+b is slope intercept form.

OpenStudy (anonymous):

i dont understand how you put them into the equation

OpenStudy (anonymous):

idk eather sorry that is all I kno...

OpenStudy (anonymous):

manipulate the variables so that y is the only thing on one side of the equation.

OpenStudy (anonymous):

y>-3x-4

OpenStudy (anonymous):

Solve this like a normal equation for y. 9x+3y≥−12. 9x+3y=−12 so you subtract 9x from both sides. 3y=-12-9x then divide by 3. y=(−12-9x)/3 then simplify. y=-12/3-9x/3 getting y=-4-3x and just rearrange it a little bit. y=-3x-4. and add back the inequality \[y \ge -3x-4\]

OpenStudy (anonymous):

9x+3y>-12 -9x -9x 3y>-9x-12 /3 /3 y>-9x/3-12/3